At what time will a train travelling at 50 kilometres per hour arrive at the stations if it is 5 kilometres from the station at 5:00 p.m
Answer by richwmiller(17219)

50*x=5
x=1/10 hour
60/10=6 minutes
it will arrive at 5:06 pm